Brianna’s age is ½ of Jessica’s age. The sum of their ages is 27 years. How old is each girl?

Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

Brianna’s age is ½ of Jessica’s age. The sum of their ages is 27 years. How old is each girl?
----------------
x + 2x = 27
etc

Answer by mananth(16946)

Brianna’s age is ½ of Jessica’s age. The sum of their ages is 27 years. How old is each girl?
Let Jessica's age be 2x
Brianna's age will be x
The sum of their ages is 27 years
2x+x=27
3x=27
divide by 3
x=9
Brianna's age will be 9 years
Let Jessica's age be 2x = 2*9 = 18 years